OpenStudy (anonymous):

whats are the factors of 9x^2=64

OpenStudy (anonymous):

a^2 - b^2 = (a+b)(a-b) 9x^2 -64 = 0 (3x)^2 - 8^2 = 0 (3x+8)(3x-8) = 0 x = 8/3 x = -8/3